# Spatial disaggregation of time series

### v1

@@ 4.2 Simulation parameters and metrics
$R^{2}$ $=1-\frac{\sum_{i=1}^{n}\sum_{t=1}^{T}(Y_{it}-\hat{Y}_{it})^{2}}{\sum_{i=1}^{n}\sum_{t=1}^{T}(Y_{it}-\bar{Y})^{2}}$
$MAPE$ $=\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}\left|\frac{Y_{it}-\hat{Y}_{it}}{Y_{it}}\right|\times 100$
$RRMSE$ $=\frac{RMSE}{\bar{Y}}=\frac{1}{\bar{Y}}\sqrt{\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}(Y_{it}-\hat{Y}_{it})^{2}}$
$RMSE$ $=\sqrt{\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}(Y_{it}-\hat{Y}_{it})^{2}}$
$\chi^{2}$ $=\sum_{i=1}^{n}\left(\frac{\sum_{t=1}^{T}(Y_{it}-\hat{Y}_{it})}{\sum_{t=1}^{T}Y_{it}}\right)$ (19)
@@
$\text{ratio}=\frac{\beta}{\sigma^{2}}\quad\text{with}\quad\begin{cases}\text{Low}&\text{if }\text{ratio}<5\\ \text{Medium}&\text{if }5\leq\text{ratio}<50\\ \text{High}&\text{if }50\leq\text{ratio}\leq 500\\ \text{Very High}&\text{if }\text{ratio}>500\end{cases}$
@@
Figure 2 shows that the behavior of the mediated RMSE and the quartiles is similar regardless of the size of the spatial grid according to the quotient between $\frac{\beta}{\sigma^{2}}$. It is clear that the higher this value, the better the reconstruction of the time series. It is also shown that as $T$ grows, there is a slight decrease in the RMSE of the disaggregation, although it is a smooth logarithmic trend.
@@
$\beta/\sigma^{2}$ n RMSE $\log$ RRMSE MAPE $R^{2}$ $\chi^{2}$
@@
Figure 2 shows that the behavior of the median RMSE and the quartiles is similar regardless of the spatial grid size based on the ratio between $\frac{\beta}{\sigma^{2}}$. It is placed on a logarithmic scale to avoid very small numbers. It is evident that the higher this value, the better the reconstruction of the time series. It also shows that, as $T$ increases, a slight decrease in the RMSE of the disaggregation is observed, although it is a smooth logarithmic trend.
The small values show that the disaggregation yields very low metric results. Table 3 shows the summary metrics for each of the ratio levels between $\frac{\beta}{\sigma^{2}}$. Since the RMSE is a non-relative measure, it increases as the ratio increases. This is explained because as the ratio increases, the variable $Y_{it}$ in the simulation takes on larger values. Therefore, by relativizing this RMSE, the RRMSE and the $\log RRMSE$ are shown to illustrate the difference between the models. This is much smaller as the ratio increases, as seen in Figure 2. The MAPE is also shown, where the reduction is even more significant, with values less than 1% for the High and Very High values. The $R^{2}$ is also shown, which obtains very good values for High and Very High. It presents medium-high $R^{2}$ values when the ratio is Medium, and negative values close to zero when the ratio is Low. This is explained because, when the variables $\boldsymbol{Z}$ are not informative in the proposed model, the disaggregation is similar to performing a disaggregation with the mean, that is, a proxy for naivety.
The value of $\chi^{2}$ proposed in the equation (19) is also presented, which weights the goodness of fit for each spatial unit. This is of particular interest because it allows us to assess whether the spatial units are well disaggregated. It is shown that the model disaggregates all variables very well for each spatial unit, even with large $n$ sizes.
@@
What was stated above regarding $R^{2}$ is reaffirmed in Figure 4. It shows that the behavior of this metric is very similar regardless of the number of spatial units. Furthermore, only in the case of a low coefficient does the model yield mediocre results, because the information provided by the covariates is too limited to reconstruct the series.
It should be noted that all the metrics shown above were calculated using the average of all $\rho$ values. This demonstrates the consistency of the method regardless of the value of the spatial correlation. However, we will now discuss the behavior of the RRMSE with respect to the spatial component of the model.
@@
Figure 4 shows the RRMSE values on a logarithmic scale. They are subdivided by the number of spatial units. Here it is shown for 9, 16, and 25. It is observed that the behavior according to the quotient $\beta/\sigma^{2}$ is similar to what we had already observed in the previous figures; that is, the higher the quotient, the better the breakdown of the time series. But this behavior only occurs for $\rho$ values of -0.5, 0.25, 0, 0.25. When $\rho$ is -0.75, very anomalous behavior occurs when n is 9. And when $\rho$ is 0.5, anomalous behavior occurs with very large $T$ values. As for the value of $\rho=0.75$, when the $T$ is greater than 36, the model is completely damaged at high ratios.
@@
In Figure 5 it is shown analogously to Figure 4 for 36, 49 and 64. It is observed that the behavior is good for all values of rhorho except when this value is 0.75 and $T$ is greater than 48. With these last two graphs it is observed that the spatial part alters the model when it is very strong, because the matrix $\mathbf{A}$ is close to being singular, and therefore, presents estimability problems. Furthermore, in a practical sense, if the variables $\boldsymbol{Z}$ do not capture the spatiality, they will make the model unstable. Therefore, this model should be applied with caution if the estimator of $\rho$ is close to 1 and the sample size in the time series is medium or large. In all other scenarios, the disaggregation is as good as the covariates that allow the disaggregation of the series of interest.
@@ 5 Application- GDP in Spain
    Industrial Production Index (IPI). The IPI is a monthly short-term indicator measuring changes in the physical volume of industrial production in Spain (excluding construction), covering sections B (Mining and quarrying), C (Manufacturing), D (Electricity and gas supply), and division 36 of section E (Water supply; sewerage, waste management, and remediation activities) of CNAE-2009. Based on a survey of over 11 500 establishments, it strips out price effects to track real output dynamics (Instituto Nacional de Estadística (2025a), INE).
@@
    Population. The annual Municipal Register provides resident counts by Autonomous Community; population size scales production and consumption, thereby influencing regional GDP levels (Instituto Nacional de Estadística (2025b), INE).
@@
    Industrial Energy Consumption. The ENC survey collects annual data on electricity, natural gas, fuel oil, renewables, petroleum products, and other fuels used by industrial firms with at least 20 employees in CNAE-2009 sections B and C. It serves as a proxy for industrial intensity (Instituto Nacional de Estadística (2022a), INE).
@@
    Business Demography. Monthly statistics on company births, dissolutions, and capital changes yield new firm counts, subscribed capital, and disbursed capital—key indicators of business dynamism and regional investment activity (Instituto Nacional de Estadística (2022b), INE).
@@
    Household Food Expenditure. Annual monetary expenditure on food by households, disaggregated by Autonomous Community, captures regional consumption patterns relevant for GDP via the expenditure approach (Ministerio de Agricultura, Pesca y Alimentación (2022), MAPA).
@@
    Labour Force and Activity Rates. Quarterly Labour Force Survey data on economically active population by sex and activity rate reflect the availability and participation of human capital across regions (Instituto Nacional de Estadística (2022), INE).
@@ 5.1 Spatial Weight Matrices
In contrast, Ceuta and Melilla consistently exhibit reddish shades in their pairwise comparisons with all other autonomous communities, highlighting their distinctiveness. These differences are likely due to their high population densities, elevated birth rates, and markedly different labor market indicators, such as unemployment. Andalusia also shows substantial distances from many other regions, suggesting a differentiated socio-economic profile characterized by higher unemployment, lower HDI, and above-average poverty risk. These patterns underscore the value of incorporating a wide range of indicators when defining spatial similarity, as doing so reveals not only geographic groupings but also deeper structural disparities across regions.
@@ 5.2 Selection of Covariates
Although some energy-related variables exhibited missing values in certain years and regions, these gaps were relatively limited in both frequency and magnitude. To preserve the internal structure and temporal consistency of each series, imputation was conducted independently for each region and variable using linear interpolation over time. This method, while simple, is widely used when the primary concern is maintaining continuity without introducing artificial trends or external information (Little and Rubin, 2002; Junger and De Leon, 2015). More sophisticated imputation strategies could potentially exploit cross-variable or cross-region correlations, but these were deliberately avoided to prevent leakage of information that could bias the temporal disaggregation model.
@@
Because many covariates exhibited strong multicollinearity—especially due to the use of multiple correlated economic indicators and energy series—a Principal Component Analysis (PCA) was applied to the standardized covariates. PCA is a classical technique for dimension reduction that transforms a correlated set of variables into a smaller number of uncorrelated components that retain most of the original variance. This transformation improves numerical stability and interpretability in regression-type models, especially when the number of predictors is large relative to the sample size or when predictors are highly collinear (Jolliffe, 2011). In the context of temporal disaggregation, the use of PCA has been shown to improve estimation by regularizing the predictor space and avoiding overfitting (Durbin, 1960; Proietti, 2006).
@@ 5.3 Results of disaggregation
$\sigma^{2}$ $0.9155$
$\phi_{1}$ $0.9983$
To assess model performance, we compared the observed and disaggregated GDP series using two metrics: Mean Absolute Percentage Error (MAPE) and Root Mean Squared Error (RMSE) relative to the mean GDP. These results are reported in Table 6. In addition to the Gower-based dissimilarity matrix, we also considered a conventional spatial contiguity matrix based on shared borders between regions. However, this alternative yielded consistently worse results in terms of both error metrics, reinforcing the advantage of the Gower approach for capturing interregional similarities in this application.
@@
$W$ Scenario MAPE Relative RMSE
@@
CCAA MAPE (%) RRMSE (%) R2 (%) Average GDP
09 3.26 4.09 94.34 202.48
13 11.03 14.55 96.81 200.97
01 7.90 9.36 95.59 144.68
10 7.54 9.59 95.61 101.33
16 4.15 4.52 97.68 63.85
12 6.65 8.37 90.09 55.69
07 7.99 9.84 73.86 53.40
05 15.02 18.04 89.58 39.27
08 3.06 4.24 95.17 38.18
02 5.87 7.51 85.92 33.52
14 7.89 12.14 89.15 28.28
04 6.08 7.43 89.09 27.60
03 18.66 26.40 84.22 21.57
11 18.06 28.50 81.85 18.22
15 17.63 22.99 61.91 18.04
06 20.52 28.45 77.24 12.52
17 20.23 27.06 40.35 7.82
Table 7: Accuracy metrics of the disaggregated GDP estimates (with anchor) for each Spanish autonomous community, identified by CCAA code. MAPE = Mean Absolute Percentage Error; RRMSE = Relative Root Mean Squared Error; $R^{2}$ = Coefficient of Determination. GDP is expressed in millions of euros.
@@
Furthermore, although the principal components capture most of the variation in auxiliary covariates, there is an implicit assumption that the relation between these covariates and the GDP is linear and stable over time. This may not hold in periods of structural economic change. Despite this limitation Figure 7 shows how the model correctly identifies the trends caused by the 2008 financial crisis and the 2020 COVID pandemic. Lastly, while the method offers strong empirical performance, further research could explore formal model selection strategies for determining the optimal number of components or penalization strength.

### v2

@@ 3.2 Simulation parameters and metrics
$R^{2}$ $=1-\frac{\sum_{i=1}^{n}\sum_{t=1}^{T}(Y_{it}-\hat{Y}_{it})^{2}}{\sum_{i=1}^{n}\sum_{t=1}^{T}(Y_{it}-\bar{Y})^{2}},$
$MAPE$ $=\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}\left|\frac{Y_{it}-\hat{Y}_{it}}{Y_{it}}\right|\cdot 100,$
$RRMSE$ $=\frac{RMSE}{\bar{Y}}=\frac{1}{\bar{Y}}\sqrt{\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}(Y_{it}-\hat{Y}_{it})^{2}},$
$RMSE$ $=\sqrt{\frac{1}{nT}\sum_{i=1}^{n}\sum_{t=1}^{T}(Y_{it}-\hat{Y}_{it})^{2}},$
$\chi^{2}$ $=\sum_{i=1}^{n}\left(\frac{\sum_{t=1}^{T}(Y_{it}-\hat{Y}_{it})}{\sum_{t=1}^{T}Y_{it}}\right).$ (26)
@@
$\text{ratio}=\frac{\beta_{1}}{\sigma^{2}}\quad\text{with}\quad\begin{cases}\text{Low}&\text{if }\text{ratio}<5,\\ \text{Medium}&\text{if }5\leq\text{ratio}<50,\\ \text{High}&\text{if }50\leq\text{ratio}\leq 500,\\ \text{Very High}&\text{if }\text{ratio}>500.\end{cases}$
@@
Figure 2 shows that the behavior of the median RRMSE and the quartiles is similar regardless of the size of the spatial grid according to the quotient between $\frac{\beta_{1}}{\sigma^{2}}$. It is clear that the higher this value, the better the reconstruction of the time series. It is also shown that as $T$ grows, there is a slight decrease in the RRMSE of the disaggregation, although it is a smooth logarithmic trend.
@@
$\beta_{1}/\sigma^{2}$ n RMSE $\log$ RRMSE MAPE $R^{2}$ $\chi^{2}$
@@
The small values show that the disaggregation yields very low metric results. Table 2 shows the summary metrics for each of the ratio levels between $\frac{\beta_{1}}{\sigma^{2}}$. Since the RMSE is a non-relative measure, it increases as the ratio increases. This is explained because as the ratio increases, the variable $Y_{it}$ in the simulation takes on larger values. Therefore, by relativizing this RMSE, the RRMSE and the $\log RRMSE$ are shown to illustrate the difference between the models. This is much smaller as the ratio increases, as seen in Figure 2. The MAPE is also shown, where the reduction is even more significant, with values less than 1% for the High and Very High values. The $R^{2}$ is also shown, which obtains very good values for High and Very High. It presents medium-high $R^{2}$ values when the ratio is Medium, and negative values close to zero when the ratio is Low. This is explained because, when the variables $\boldsymbol{Z}$ are not informative in the proposed model, the disaggregation is similar to performing a disaggregation with the mean, that is, a proxy for naivety (See $\hat{\mathbf{Y}}^{1}$ estimator in Section 2.3).
The value of $\chi^{2}$ proposed on the Equation (26) weights the goodness of fit for each spatial unit. This is of particular interest because it allows us to assess whether the spatial units are well disaggregated. It is shown that the model disaggregates all variables very well for each spatial unit, even with large $n$ sizes.
@@
What was stated above regarding $R^{2}$ is reaffirmed in Figure 3. In Table 2, it is shown that the behavior of this metric is very similar regardless of the value of $n$. Furthermore, only in the case of a low coefficient $\frac{\beta_{1}}{\sigma^{2}}$ does the model yield mediocre results, because the information provided by the covariates is too limited to reconstruct the series.
It should be noted that all the metrics shown above were calculated using the average of all $\rho$ values. This demonstrates the consistency of the method regardless of the value of the spatial correlation. However, we will now discuss the behavior of the RRMSE with respect to $\rho$, the spatial component of the model.
@@
Figure 4 shows the mean RRMSE values on a logarithmic scale. They are subdivided by the number of spatial units. Here it is shown for 9, 16, and 25. It is observed that the behavior according to the quotient $\beta/\sigma^{2}$ is similar to what we had already observed in the previous figures; that is, the higher the quotient, the better the breakdown of the time series. But this behavior only occurs for $\rho$ values of -0.5, 0.25, 0, and 0.25. When $\rho$ is $\pm 0.75$, very anomalous behavior occurs when $n$ is 9. And when $\rho$ is 0.5, anomalous behavior occurs with very large $T$ values. As for the value of $\rho=0.75$, when the $T$ is greater than 36, the model is completely damaged at high ratios. This is justified by the results obtained in Theorem 2.
@@
In Figure 5 it is shown analogously to Figure 4 for 36, 49, and 64. It is observed that the behavior is good for all values of $\rho$ except when this value is 0.75 and $T$ is greater than 48. With these last two figures, it is observed that the spatial part alters the model when it is very strong, because the matrix $\mathbf{A}$ is close to being singular, and therefore, presents estimability problems. Furthermore, in a practical sense, if the variables $\boldsymbol{Z}$ do not capture the spatiality, they will make the model unstable. Therefore, this model should be applied with caution if the estimator of $\rho$ is close to 1 and the sample size in the time series is medium or large. In all other scenarios, the disaggregation is as good as the covariates that allow the disaggregation of the series of interest.
@@ 4 Application- GDP in Spain
    Industrial Production Index (IPI). The IPI is a monthly short-term indicator measuring changes in the physical volume of industrial production in Spain (excluding construction), covering sections B (Mining and quarrying), C (Manufacturing), D (Electricity and gas supply), and division 36 of section E (Water supply; sewerage, waste management, and remediation activities) of CNAE-2009. Based on a survey of over 11 500 establishments, it strips out price effects to track real output dynamics Instituto Nacional de Estadística (2025a) (INE).
@@
    Population. The annual Municipal Register provides resident counts by Autonomous Community; population size scales production and consumption, thereby influencing regional GDP levels Instituto Nacional de Estadística (2025b) (INE).
@@
    Industrial Energy Consumption. The ENC survey collects annual data on electricity, natural gas, fuel oil, renewables, petroleum products, and other fuels used by industrial firms with at least 20 employees in CNAE-2009 sections B and C. It serves as a proxy for industrial intensity Instituto Nacional de Estadística (2022a) (INE).
@@
    Business Demography. Monthly statistics on company births, dissolutions, and capital changes yield new firm counts, subscribed capital, and disbursed capital—key indicators of business dynamism and regional investment activity Instituto Nacional de Estadística (2022b) (INE).
@@
    Household Food Expenditure. Annual monetary expenditure on food by households, disaggregated by Autonomous Community, captures regional consumption patterns relevant for GDP via the expenditure approach Ministerio de Agricultura, Pesca y Alimentación (2022) (MAPA).
@@
    Labour Force and Activity Rates. Quarterly Labour Force Survey data on the economically active population by sex and activity rate reflect the availability and participation of human capital across regions Instituto Nacional de Estadística (2022) (INE).
@@ 4.1 Spatial Weight Matrices
In contrast, Ceuta and Melilla consistently exhibit reddish shades in their pairwise comparisons with all other autonomous communities, highlighting their distinctiveness. These differences are likely due to their high population densities, elevated birth rates, and markedly different labor market indicators, such as unemployment. Andalucía also shows substantial distances from many other regions, suggesting a differentiated socio-economic profile characterized by higher unemployment, lower HDI, and above-average poverty risk. These patterns underscore the value of incorporating a wide range of indicators when defining spatial similarity, as doing so reveals not only geographic groupings but also deeper structural disparities across regions.
@@ 4.2 Selection of Covariates
Although some energy-related variables exhibited missing values in certain years and regions, these gaps were relatively limited in both frequency and magnitude. To preserve the internal structure and temporal consistency of each series, imputation was conducted independently for each region and variable using linear interpolation over time. This method, while simple, is widely used when the primary concern is maintaining continuity without introducing artificial trends or external information Little and Rubin (2002); Junger and De Leon (2015). More sophisticated imputation strategies could potentially exploit cross-variable or cross-region correlations, but these were deliberately avoided to prevent leakage of information that could bias the temporal disaggregation model.
@@
Because many covariates exhibited strong multicollinearity—especially due to the use of multiple correlated economic indicators and energy series—a Principal Component Analysis (PCA) was applied to the standardized covariates. PCA is a classical technique for dimension reduction that transforms a correlated set of variables into a smaller number of uncorrelated components that retain most of the original variance. This transformation improves numerical stability and interpretability in regression-type models, especially when the number of predictors is large relative to the sample size or when predictors are highly collinear Jolliffe (2011). In the context of temporal disaggregation, the use of PCA has been shown to improve estimation by regularizing the predictor space and avoiding overfitting Durbin (1960); Proietti (2006).
@@ 4.3 Results of disaggregation
$\sigma^{2}$ $0.2707$
$\phi_{1}$ $0.8961$
To assess model performance, we compared the observed and disaggregated GDP series using two metrics: Mean Absolute Percentage Error (MAPE) and Relative Root Mean Squared Error (RRMSE) relative to the mean GDP. These results are reported in Table 5. In addition to the Gower-based dissimilarity matrix, we also considered a conventional spatial contiguity matrix based on shared borders between regions. However, this alternative yielded consistently worse results in terms of both error metrics, reinforcing the advantage of the Gower approach for capturing interregional similarities in this application.
@@
$W$ Scenario MAPE RRMSE
@@
CCAA MAPE (%) RRMSE (%) R2 (%) GDP N-MAPE(%)
09 3.26 4.09 94.34 202.48 68.98
13 11.03 14.55 96.81 200.97 68.56
01 7.90 9.36 95.59 144.68 56.64
10 7.54 9.59 95.61 101.33 38.16
16 4.15 4.52 97.68 63.85 1.83
12 6.65 8.37 90.09 55.69 12.80
07 7.99 9.84 73.86 53.40 17.02
05 15.02 18.04 89.58 39.27 59.43
08 3.06 4.24 95.17 38.18 64.67
02 5.87 7.51 85.92 33.52 87.26
14 7.89 12.14 89.15 28.28 122.82
04 6.08 7.43 89.09 27.60 128.94
03 18.66 26.40 84.22 21.57 189.90
11 18.06 28.50 81.85 18.22 244.69
15 17.63 22.99 61.91 18.04 247.72
06 20.52 28.45 77.24 12.52 399.75
17 20.23 27.06 40.35 7.82 700.91
Table 6: Accuracy metrics of the disaggregated GDP estimates (with anchor) for each Spanish autonomous community, identified by CCAA code. MAPE = Mean Absolute Percentage Error; RRMSE = Relative Root Mean Squared Error; $R^{2}$ = Coefficient of Determination. GDP mean is expressed in millions of euros, N-MAPE = Mean Absolute Percentage Error of the Naive Estimator
@@
In addition, we computed the mean absolute percentage error (N-MAPE) of the naive estimator proposed in Equation (19). The regional MAPEs reveal a poor adjustment to the observed data, with values ranging from $1.83\%$ to $700.91\%$, and a median above $60\%$. In particular, several regions display extremely large discrepancies, indicating that the block-averaging approach fails to reproduce the true disaggregated dynamics. This poor performance is consistent with the theoretical results derived in Theorem 2: when the conditions that make the naive and the proposed estimators equivalent are not simultaneously satisfied—specifically, when $\rho$ departs from 0 and the regressors $\mathbf{Z}$ exhibit marked spatial heterogeneity—the proposed estimator achieves a substantially lower mean squared error than naive estimator.
The numerical results from the applied example strongly support the superiority of the proposed disaggregation estimator over the naive block-average predictor. In particular, the decomposition of $\Delta$ in Theorem 2 yields the following contributions (computed at the estimated parameter values): $\boldsymbol{\mu}^{\top}(\mathbb{I}_{nT}-\mathbf{P})\boldsymbol{\mu}\approx 2.58\times 10^{4},\quad 2k_{\rho}\mathbf{m}^{\top}\mathbf{m}\approx 5.00,\quad 2\boldsymbol{\mu}^{\top}(\mathbb{I}_{nT}-\mathbf{P})(\mathbb{I}_{T}\otimes\mathbf{D}_{\rho})(\mathbf{Y}_{a}-\boldsymbol{\mu}_{a})\approx 42.8,$ while the parameter-uncertainty contribution is $\operatorname{tr}(\mathbf{M}\operatorname{Var}(\hat{\beta})\mathbf{M}^{\top})\approx 273.4$. The first term, which measures the spatial heterogeneity of the fitted systematic component $\boldsymbol{\mu}=\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}$ not captured by block averaging, overwhelmingly dominates the decomposition. Consequently, $\Delta$ is large and positive, implying a substantial reduction in MSE when using the spatially informed predictor.
Note that, in the example, only one of the five conditions, high temporal persistence, $|\phi_{1}|\approx 1$) holds (approximately); the others do not: $\boldsymbol{\beta}^{\top}\boldsymbol{\beta}$ is not small, $\mathbf{Z}$ displays nontrivial within-period spatial variation ($\|\mathbf{Z}-\frac{1}{n}\mathbf{C}^{\top}\mathbf{CZ}\|\approx 1967$), and $\rho$ is moderate ($\rho\approx 0.614$), so the large gain is driven primarily by spatial heterogeneity in $\mathbf{A}^{-1}\mathbf{Z}\boldsymbol{\beta}$.
@@
Furthermore, although the principal components capture most of the variation in auxiliary covariates, there is an implicit assumption that the relation between these covariates and the GDP is linear and stable over time. This may not hold in periods of structural economic change. Despite this limitation, Figures 7 and 8 show how the model correctly identifies the trends caused by the 2008 financial crisis and the 2020 COVID pandemic. Lastly, while the method offers strong empirical performance, further research could explore formal model selection strategies for determining the optimal number of components or penalization strength.